Solve: x3 +6=2x2 + 5x use factor and remainder theorem
Answer #1

solution given that x + 6 =. 5x = 0 x² + 6 2x²+5x 226² x² - 2x² - 5x + 6 U let P. (2) Putting r =) - 5+ 6 7 Pl= 7 - 7 po rema